# Estimating the Lateral Motion States of an Underwater Robot by Propeller Wake Sensing Using an Artificial Lateral Line

### v1

@@ Abstract
An artificial lateral line (ALL) is a bioinspired flow sensing system of an underwater robot that consists of distributed flow sensors. The ALL has achieved great success in sensing the motion states of bioinspired underwater robots, e.g., robotic fish, that are driven by body undulation and/or tail flapping. However, the ALL has not been systematically tested and studied in the sensing of underwater robots driven by rotating propellers due to the highly dynamic and complex flow field therein. This paper makes a bold hypothesis that the distributed flow measurements sampled from the propeller wake flow, although infeasible to represent the entire flow dynamics, provides sufficient information for estimating the lateral motion states of the leader underwater robot. An experimental testbed is constructed to investigate the feasibility of such a state estimator which comprises a cylindrical ALL sensory system, a rotating leader propeller, and a water tank with a planar sliding guide. Specifically, a hybrid network that consists of a one-dimensional convolution network (1DCNN) and a bidirectional long short-term memory network (BiLSTM) is designed to extract the spatiotemporal features of the time series of distributed pressure measurements. A multi-output deep learning network is adopted to estimate the lateral motion states of the leader propeller. In addition, the state estimator is optimized using the whale optimization algorithm (WOA) considering the comprehensive estimation performance. Extensive experiments are conducted the results of which validate the proposed data-driven algorithm in estimating the motion states of the leader underwater robot by propeller wake sensing.
@@ I Introduction
In the nature, the lateral line is an essential class of fish’s flow-sensing organ that plays a crucial role in their flow-relative behaviors such as obstacle avoidance, rheotaxis, predation, and schooling[1, 2]. Consisting of superficial and canalicular neuromasts, the lateral line measures the fluid vibration and the pressure gradient to perceive the surrounding flow environment[3]. Taking the inspiration therein, researchers have designed various artificial lateral line systems (ALLs) aiming to enhance the flow-sensing capability of bioinspired underwater robots[4, 5, 6, 7, 8, 9, 10, 11].
Most existing research on the ALL has focused on the sensing of flow fields generated by either bioinspired robotic fish with body/fin undulations or dipoles with periodic oscillations. Generally speaking, such flow fields are tractable and represented by either analytical potential flow models[12, 13] or regular vortex street models[14, 15]. Based on these hidden flow models, the sensory measurements of the ALL are sufficient to capture the main coherent flow structures, thus leading to successful estimation of the flow field of interest. However, the majority of underwater robots used in the real world are equipped with high-speed rotating propellers[16, 17, 18], whose wake flows are highly dynamic and complex compared to those of bioinspired fish robots[19, 20, 21, 22, 23]. The modeling and analysis of the propeller wake typically requires high-precision and high-sensitivity computational fluid dynamics simulation, which is computationally expensive inhibiting real-time flow sensing[24, 25]. The dynamic and complex flow field of the propeller wake presents a significant challenge for the application of the ALL therein. Therefore, although many successful designs and applications of ALLs have been demonstrated, there still remains an open question whether ALLs are capable of providing sufficient feedback information for flow sensing of propeller-driven underwater robots, and if yes, how to design an estimation algorithm to assimilate the distributed measurements.
@@
The contributions of this paper are twofold. First, this paper presents, to the best of the authors’ knowledge, the first efforts to investigate the feasibility and the algorithmic design of flow sensing with highly dynamic and complex propeller wake using an ALL in a leader-follower formation of underwater robots. A testbed is designed and extensive experiments are conducted to provide insights into this challenging problem. Second, this paper proposes a data-driven algorithm for estimating the lateral motion states of the propeller-driven leader robot using the ALL. The algorithm integrates 1DCNN and BiLSTM to extract spatiotemporal flow features as well as a multi-output deep neutral network for the lateral motion state estimation. In addition, to improve the accuracy of the estimation results, the task weights of the loss function designed for the multi-state estimation are optimized using the bioinspired WOA algorithm.
@@ II Problem Description
Figure 1: Illustration of the ALL sensing the propeller wake of a leader underwater robot in a leader-follower formation (Top View).
Leader-follower formation is one of the most commonly-used formation configuration in underwater robots[26]. Analogous to the lateral line of biological fish in sensing a leader fish, the ALL is expected to estimate the corresponding motion states of the leader robot by sampling its wake flow (Fig. 1). Particularly, we are more interested in sensing the lateral motion of the leader compared to the longitudinal motion which has a significant influence on the formation control. Unlike the robotic fish whose wake flow is typically generated by the body undulation and/or fin flapping, underwater robots driven by propellers have much more complex wake structures coming from the interaction of the high-speed rotating propellers and the surrounding fluid. In this paper, we hypothesize that the spatiotemporal flow measurement data acquired by the ALL, although only contains partial coherent flow information of the highly dynamic propeller wake, holds sufficient information regarding the lateral motion of the leader robot, which is described by the velocity magnitude, the velocity direction and the displacement.
@@ II-A Testing Platform
The wake of a propeller-driven underwater robot travelling through aquatic environments is mainly dominated by the propeller wake, therefore, this paper simplifies the problem and focuses on the estimation of the motion states of a leading propeller. For this purpose, we have constructed a testing apparatus shown in Fig. 2. This experimental testbed mainly consists of a testing water tank, a high-precision sliding guide, a data acquisition system and a leader propeller. The water tank is 3m long, 2m wide and 1.5m deep. The sliding guide provides support and high-precision movement between the rotating propeller and the ALL system. A high-performance industrial personal computer (IPC) controls the sliding guide with a linear travelling velocity between 0.2 m/min and 1.2 m/min and a positional accuracy of 1 mm. Similarly to a towing tank, the sliding guide is rigidly attached to the water tank and is controllable through preprogramming or manual operation.
Figure 2: The designed testing platform that comprises a water tank, a sliding guide, a leader propeller, and an ALL.
Attached to the sliding guide are a data acquisition system and a leader propeller system. The data acquisition system comprises an ALL equipped with three pressure sensors and a microcomputer. A circular shaped cross-section is widely used in ALL flow sensing for its corresponding tractable hydrodynamic properties. This paper adopts the same circular shape design for the ALL (Fig. 2). The cylindrical ALL has a radius of 4 cm and a height of 20 cm. Three through-holes with a diameter of 4 mm are evenly placed for holding pressure sensors with an angle of 45 degrees between each adjacent pair at the same horizontal level of 10 cm from the bottom of the cylindar. The pressure sensors used are MS5837-02Ba, featuring a resolution of 1 Pa and I2C bus communication. To acquire multiple sensors simultaneously in real time, an TI TCA9548 eight-channel bidirectional transfer switch is used.
@@ III-A Data Acquisition
We conduct extensive experiments using the aforementioned testbed to create a dataset of the ALL measurements given selected lateral motion states of the leader propeller. To eliminate the influence of the submerged depth of the pressure sensors, all the sensor outputs are debiased by deducting the corresponding pressure readings sampled before the rotation of the propeller when the flow is stationary. In each experimental trial, the leader propeller moves along the lateral axis or the $x$-axis at a constant speed $v$ with motion direction $d\in\{N,P\}$ while the longitudinal displacement $y$ is kept constant. We select five representative speeds of the moving propeller, i.e., $v\in\{400,500,600,700,800\}$ mm/s. In each trial, the sliding guide accelerates the propeller from at rest to the selected speed $v$ and keeps it at that constant speed. Considering the effective measurable width of the propeller wake by the ALL and the influence of the transient during the propeller accelerating period, the lateral motion of the propeller goes from $x=-175$ mm to $x=175$ mm and then back to $x=-175$ mm. The effective estimation range is clipped to $x\in[-120,120]$ mm and normalized in the following sections for the convenience of presentation. In addition, two longitudinal displacements are selected in the experiment, i.e., $y\in\{250,300\}$ mm.
@@
Figure 3 shows the collected distributed pressure data given selected speeds and directions of the leader propeller. We observe strong correlations between time-domain features of the pressure measurements and the lateral motion states of interest. Specifically, the changing trends of the collective pressure measurements are primarily correlated to the direction of motion of the leader propeller; the changing rates of the pressure measurements are mainly correlated to the traveling speed of the propeller; and the magnitudes of the pressure measurements are mainly correlated to the lateral displacement of the propeller. Such observations indicate that the ALL pressure measurements of the propeller wake (most likely) contains sufficiently rich information for the estimation of the lateral motion states of the leader propeller, supporting our previous hypothesis.
@@ III-B Time Series Selection
Most existing flow sensing studies using ALLs deal with structured flow fields such as the Von Karman vortex street of laminar flow passing a cylinder or periodic wake flow of an undulating fish-like fin. This paper, however, targets the propeller wake which features cross-scale dynamic and unstable flow structures, thus making the flow sensing problem extremely challenging. Due to the practical limitation of sampling frequency and sampling resolution of the All, it is infeasible to accurately capture all the characteristics of the propeller wake. Additionally, the phase averaging approach commonly adopted in propeller wake analysis requires the data sampling to occur at uniform phases of the periodic rotation of the propeller, which exceeds the sampling capability of almost all the ALL sensing systems. Through high-fidelity CFD simulation, we find that during the motion of the propeller, there exist a stochastic pulsation process in the ALL pressure output along with statistically distinguishable trends corresponding to different motion configurations. Therefore, we hypothesize that the time series of ALL measurement data provides partial but sufficient information regarding the lateral motion of the propeller of interest. Furthermore, increasing the number of sensors and/or the total sampling time leads to added information acquired, enhancing the robustness of the flow sensing algorithm, which poses a balancing problem between the practical engineering considerations (e.g., the cost, the installation space and real-time calculation) and the estimation accuracy. Specifically, we define the time series of the $i$-th pressure sensor output at time $k$ used for wake flow sensing as
@@ IV-A Flow Feature Extraction with Hybrid CNN-BiLSTM
Considering the highly complex and stochastic dynamics of propeller wake, we adopt the trending powerful deep learning approach to extract hidden spatiotemporal flow features given the time series of the distributed ALL measurements and design a hybrid function of the one-dimension convolutional neural network (1DCNN) and the bidirectional long short-term memeory network (BiLSTM). Specifically, as one of the most widely used networks[27, 28], 1DCNN is selected to extract hidden spatial features of the sequential pressure measurement data supported by sparse connections and weight sharing. The network design consists of four main layers including the input, the convolutional, the batch normalization, and the pooling (Fig. 4). On the other hand, BiLSTM is a special recurrent neural network (RNN)[29, 30] which introduces a gate mechanism to control both the forward and reverse information transmission paths, rendering BiLSTM a stronger learning ability for long-range dependency compared to traditional RNNs[29]. In this paper, the BiLSTM network follows 1DCNN and learns the bidirectional sequential relationship from the extracted spatial features, thereby providing long-term characteristics of the time series of ALL data sampled from the unsteady and complex wake flow of the leader propeller.
@@ IV-B Motion State Estimator
Design of the loss function. This paper focuses on the estimation of the lateral motion states of the leader propeller including the relative displacement $x$, travelling speed $v$, and direction of motion $d$. Considering the properties of the motion states, we model the estimation task as a combination of regression and classification problem with continuous estimation state space of displacement $\mathcal{X}$ and discretized estimation state space of travelling speed $\mathcal{V}$ and Boolean-type direction $\mathcal{D}$. Correspondingly, we design a multi-output neural network as the motion state estimator that takes the wake flow features as input and estimates the lateral motion states of the leader propeller utilizing parameter sharing [33]. The network architecture of the designed multi-output model is depicted in the right column of Fig. 4. Given sufficiently rich experimental data, the key to the design of the state estimation model lies in the design of the loss function [34]. Considering the continuous and discrete motion states to be estimated, we first design three individual loss functions for the three estimation states, respectively. Specifically, the relative lateral displacement $x$ is estimated by regression, with its loss function $l_{1}$ designed as the mean square error loss (MSE), i.e.,
@@
Optimizing task weights. Considered as a multi-objective optimization problem, the selection of the weights $\lambda_{i}$ where $i\in\{1,2,3\}$ greatly affects the overall estimation performance. In this paper, a nature inspired meta-heuristic optimization algorithm named the whale optimization algorithm (WOA) is used to search for the optimal weights $\lambda_{i}$. Proposed by Mirjalili et al. in 2016[35], WOA mimics the hunting behavior of humpback whales in their bubble-net predation, featuring the advantage of less parameters to tune and easiness to implement. The primary process of the algorithm consists of three essential stages including encircling prey, feeding with bubble nets, and finding prey[36]. To optimize the task weights, we design a performance evaluation vector $\boldsymbol{e}=[1-\text{ACC}_{1},1-\text{ACC}_{2},\text{ERR}]$ where $\text{ACC}_{1}$ and $\text{ACC}_{2}$ represent the state estimation accuracy of the lateral motion speed $v$ and the direction of motion $d$, respectively, and ERR represents the mean square error of the estimated displacement $x$. The optimization fitness $g$ (or the cost) is then designed as 1-norm of the performance evaluation vector $\boldsymbol{e}$, i.e.,
@@ IV-C Hyperparameters and Training Environment
Figure 5: The experimental results of the motion state estimation based on propeller wake sensing including specifically the regression results on the lateral displacement $x$, and the classification results on the traveling speed $v$ and the direction of motion $d$.

### v2

@@ Abstract
The artificial lateral line (ALL), comprising distributed flow sensors, has been successful in sensing motion states of bioinspired underwater robots like robotic fish. However, its application to robots driven by rotating propellers remains unexplored due to the complexity of propeller wake flow. This paper investigates the feasibility of using ALL to sense propeller wake for underwater robot leader-follower formation. To estimate the lateral motion states of a leader propeller, this paper designs a multi-output deep learning network that extracts temporal and spatial features from distributed pressure measurements of propeller wake. Extensive experiments are conducted on a designed testbed, the results of which validate the effectiveness of the proposed propeller wake sensing method.
@@ I Introduction
In the nature, the lateral line is an essential class of fish’s flow-sensing organ that plays a crucial role in their flow-relative behaviors such as obstacle avoidance, rheotaxis, predation, and schooling[1, 2, 3]. Taking the inspiration therein, researchers have designed various artificial lateral line systems (ALLs) aiming to enhance flow-sensing capabilities of bioinspired underwater robots[4, 5, 6, 7, 8, 9, 10, 11]. Most existing research on the ALL has focused on sensing flow fields generated by bioinspired robotic fish with body/fin undulations or dipoles with periodic oscillations. Generally speaking, such flow fields are tractable and represented by either analytical potential flow models[12, 13] or regular vortex street models[14, 15].
However, most underwater robots used in the real world use propellers for propulsion [16, 17, 18]. The wake of high-speed rotating propellers is complex, dynamic and even stochastic [19, 20, 21]. Conventional instantaneous pressure measurements are insufficient in providing usable information for flow sensing, while only flow features extracted from the pressure measurement time series are potentially useful. Additionally, analyzing propeller wake demands high-precision and high-frequency sampling [22, 23, 24], which is costly and impractical for real-time robot sensing. The distributed low-cost pressure sensors, adopted in ALLs, capture limited dynamics information of the propeller wake at a frequency much lower than the rotating frequency of the propeller. This complexity challenges traditional estimation methods especially the commonly-adopted frequency analysis-based approach [25, 26, 27, 28]. Therefore, although many successful designs and applications of ALL have been demonstrated, there remains an open question of whether ALL can provide sufficient feedback information for flow sensing of propeller-driven underwater robots, and if yes, how to design an estimation algorithm to assimilate the distributed measurements.
@@
The main contributions of this paper are twofold. Firstly, by leveraging an artificial lateral line (ALL) composed of distributed commercial low-cost pressure sensors, this study explores a novel realm of sensing the highly dynamic and complex propeller wake to estimate the position of a leader high-speed rotating propeller. The wake flow under investigation is encountered more commonly by existing underwater vehicles and differs inherently from the flow fields studied in other ALL flow sensing research, such as those generated by bioinspired body undulation and/or tail flapping. Secondly, this paper proposes a novel data-driven flow sensing algorithm for estimating the lateral motion states of a leader propeller. This algorithm combines deep learning with a biomimetic optimizer to address the complex and dynamic propeller wake sensing.
@@ II Problem Description
Figure 1: Illustration of ALL sensing the propeller wake of a leader underwater robot in a leader-follower formation (Top View).
Leader-follower formation is one of the most commonly-used formation configuration in underwater robots[29]. Analogous to the lateral line of biological fish in sensing a leader fish, the ALL is expected to estimate the corresponding motion states of the leader robot by sampling its wake flow (Fig. 1). Particularly, we are more interested in sensing the lateral motion of the leader compared to the longitudinal motion which has a significant influence on the formation control. In addition, this paper replaces the leader robot with a single propeller with an aim to focus on the fundamental research problem in propeller wake sensing. While we acknowledge that wake flows generated by traveling robot bodies affect those produced by propellers, many existing studies[30, 26, 31] on ALL-based motion state estimation adopted the same simplification measures. The underlying reason is that wake flows generated by propulsion systems typically possess significantly different features than flow fields generated by traveling robot bodies, therefore, are detectable when mixed with vortices shed by traveling robots. In this paper, we hypothesize that the spatiotemporal flow measurement data acquired by the ALL, although only contains partial coherent flow information of the highly dynamic propeller wake, holds sufficient information regarding the lateral motion of the leader robot, which is described by the velocity magnitude, the velocity direction and the displacement.
@@ II-A Testing Platform
The wake of a propeller-driven underwater robot traveling through aquatic environments is mainly dominated by the propeller wake, therefore, this paper simplifies the problem and focuses on the estimation of the motion states of a leading propeller. For this purpose, we have constructed a testing apparatus shown in Fig. 2. This experimental testbed mainly consists of a testing water tank, a high-precision sliding guide, a data acquisition system and a leader propeller. The water tank is 3m long, 2m wide and 1.5m deep. The sliding guide provides support and high-precision movement between the rotating propeller and the ALL system. A high-performance industrial personal computer (IPC) controls the sliding guide with a linear traveling velocity between 0.2 m/min and 1.2 m/min and a positional accuracy of 1 mm. Similarly to a towing tank, the sliding guide is rigidly attached to the water tank and is controllable through preprogramming or manual operation.
(a) CAD design of the testing platform.
(b) ALL design with pressure sensors.
Figure 2: The design of the testing platform that comprises a water tank, a sliding guide, a leader propeller, and an ALL. The water tank measures 3.0m(L) x 2.0m(W) x 1.5m(H). The ALL consists of three pressure sensors distributively positioned with an angle of $45^{\circ}$ between each adjacent pair at the same horizontal level from the bottom of the cylindrical shell.
Attached to the sliding guide are a data acquisition system and a leader propeller system. The data acquisition system comprises an ALL equipped with three pressure sensors and a microcomputer. A circular shaped cross-section is widely used in ALL flow sensing for its corresponding tractable hydrodynamic properties. This paper adopts the same circular shape design for the ALL (as shown in Fig. 2(b)). The cylindrical ALL has a diameter of 8 cm and a height of 20 cm. Three through-holes with a radius of 2 mm are evenly placed for holding pressure sensors with an angle of 45 degrees between each adjacent pair at the same horizontal level of 10 cm from the bottom of the cylinder. The pressure sensors used are MS5837-02Ba, featuring a resolution of 1 Pa and I2C bus communication. To acquire multiple sensors simultaneously in real time, an TI TCA9548 eight-channel bidirectional transfer switch is used. Moreover, the two subfigures of Fig. 2(a) show the relative position relationship between ALL and the propeller.
@@ III-A Data Acquisition
We conduct extensive experiments using the aforementioned testbed to create a dataset of the ALL measurements given selected lateral motion states of the leader propeller. To eliminate the influence of the submerged depth of the pressure sensors, all the sensor outputs are debiased by deducting the corresponding pressure readings sampled before the rotation of the propeller when the flow is stationary. In each experimental trial, the leader propeller moves along the lateral axis or the $x$-axis at a constant speed $v$ with motion direction $d\in\{N,P\}$ while the longitudinal displacement $y$ is kept constant. We select five representative speeds of the moving propeller, i.e., $v\in\{400,500,600,700,800\}$ mm/s. Considering the effective measurable width of the propeller wake by the ALL and the influence of the transient during the propeller accelerating period, the lateral motion of the propeller goes from $x=-175$ mm to $x=175$ mm and then back to $x=-175$ mm. The effective estimation range is clipped to $x\in[-120,120]$ mm and normalized in the following sections for the convenience of presentation. In addition, two longitudinal displacements are selected in the experiment, i.e., $y\in\{250,300\}$ mm.
@@
Figure 3 shows the collected distributed pressure data given selected speeds and directions of the leader propeller. We observe strong correlations between time-domain features of the pressure measurements and the lateral motion states of interest. Specifically, the changing trends of the collective pressure measurements are primarily correlated to the direction of motion of the leader propeller; the changing rates of the pressure measurements are mainly correlated to the traveling speed of the propeller; and the magnitudes of the pressure measurements are mainly correlated to the lateral displacement of the propeller. Such observations indicate that the ALL pressure measurements of the propeller wake (most likely) contain sufficiently rich information for the estimation of the lateral motion states of the leader propeller, supporting our previous hypothesis.
@@ III-B Time Series Selection
Most existing flow sensing studies using ALLs deal with structured flow fields such as the Von Karman vortex street of laminar flow passing a cylinder or periodic wake flow of an undulating fish-like fin. This paper, however, targets the propeller wake which features cross-scale dynamic and unstable flow structures, thus making the flow sensing problem extremely challenging. Due to the practical limitation of sampling frequency and sampling resolution of the ALL, it is infeasible to accurately capture all the characteristics of the propeller wake. Additionally, the phase averaging approach commonly adopted in propeller wake analysis requires the data sampling to occur at uniform phases of the periodic rotation of the propeller, which exceeds the sampling capability of almost all the ALL sensing systems. Through high-fidelity CFD simulation, we find that during the motion of the propeller, there exists a stochastic pulsation process in the ALL pressure output along with statistically distinguishable trends corresponding to different motion configurations. Therefore, we hypothesize that the time series of ALL measurement data provides partial but sufficient information regarding the lateral motion of the propeller of interest. Furthermore, increasing the number of sensors and/or the total sampling time leads to added information acquired, enhancing the robustness of the flow sensing algorithm, which poses a balancing problem between the practical engineering considerations (e.g., the cost, the installation space and real-time calculation) and the estimation accuracy. Specifically, we define the time series of the $i$-th pressure sensor output at time $k$ used for wake flow sensing as
@@ IV-A Flow Feature Extraction with Hybrid CNN-BiLSTM
Considering the highly complex and stochastic dynamics of propeller wake, we adopt the trending powerful deep learning approach to extract hidden spatiotemporal flow features given the time series of the distributed ALL measurements and design a hybrid function of the one-dimension convolutional neural network (1DCNN) and the bidirectional long short-term memory network (BiLSTM). Specifically, as one of the most widely used networks[32, 33], 1DCNN is selected to extract hidden spatial features of the sequential pressure measurement data supported by sparse connections and weight sharing. The network design consists of four main layers including the input, the convolutional, the batch normalization, and the pooling (Fig. 4). On the other hand, BiLSTM is a special recurrent neural network (RNN)[34, 35] which introduces a gate mechanism to control both the forward and reverse information transmission paths, rendering BiLSTM a stronger learning ability for long-range dependency compared to traditional RNNs[34]. In this paper, the BiLSTM network follows 1DCNN and learns the bidirectional sequential relationship from the extracted spatial features, thereby providing long-term characteristics of the time series of ALL data sampled from the unsteady and complex wake flow of the leader propeller.
@@ IV-B Motion State Estimator
Design of the loss function. This paper focuses on the estimation of the lateral motion states of the leader propeller including the relative displacement $x$, traveling speed $v$, and direction of motion $d$. Considering the properties of the motion states, we model the estimation task as a combination of regression and classification problems with continuous estimation state space of displacement $\mathcal{X}$ and discretized estimation state space of traveling speed $\mathcal{V}$ and Boolean-type direction $\mathcal{D}$. Correspondingly, we design a multi-output neural network as the motion state estimator that takes the wake flow features as input and estimates the lateral motion states of the leader propeller utilizing parameter sharing [38]. The network architecture of the designed multi-output model is depicted in the right column of Fig. 4. Given sufficiently rich experimental data, the key to the design of the state estimation model lies in the design of the loss function [39]. Considering the continuous and discrete motion states to be estimated, we first design three individual loss functions for the three estimation states, respectively. Specifically, the relative lateral displacement $x$ is estimated by regression, with its loss function $l_{1}$ designed as the mean square error loss (MSE), i.e.,
@@
Optimizing task weights. To ensure a balanced influence of the task weights on the estimation performance, we formulate weight selection as a multi-objective optimization problem. Considering the network complexity and extended computational time, we choose the Whale Optimization Algorithm (WOA) due to its effectiveness in quickly finding optimal solutions. Inspired by humpback whales’ predatory behaviors, WOA offers advantages like minimal adjustable parameters, straightforward operation, and a robust ability to escape local optima [40, 41, 42, 43, 44, 45]. To optimize the task weights, we design a performance evaluation vector $\boldsymbol{e}=[1-\text{ACC}_{1},1-\text{ACC}_{2},\text{ERR}]$ where $\text{ACC}_{1}$ and $\text{ACC}_{2}$ represent the state estimation accuracy of the lateral motion speed $v$ and the direction of motion $d$, respectively, and ERR represents the mean square error of the estimated displacement $x$. The optimization fitness $g$ (or the cost) is then designed as 1-norm of the performance evaluation vector $\boldsymbol{e}$, i.e.,
@@ IV-C Hyperparameters and Training Environment
Figure 5: The experimental results of the motion state estimation based on propeller wake sensing including the regression results on the lateral displacement $x$, and the classification results on the traveling speed $v$ and the direction of motion $d$.